# Bi-Mamba4TS: Bidirectional Mamba for Time Series Forecasting

### v1

@@ Abstract
Long-term time series forecasting (LTSF) provides longer insights into future trends and patterns. In recent years, deep learning models especially Transformers have achieved advanced performance in LTSF tasks. However, the quadratic complexity of Transformers rises the challenge of balancing computaional efficiency and predicting performance. Recently, a new state space model (SSM) named Mamba is proposed. With the selective capability on input data and the hardware-aware parallel computing algorithm, Mamba can well capture long-term dependencies while maintaining linear computational complexity. Mamba has shown great ability for long sequence modeling and is a potential competitor to Transformer-based models in LTSF. In this paper, we propose Bi-Mamba4TS, a bidirectional Mamba for time series forecasting. To address the sparsity of time series semantics, we adopt the patching technique to enrich the local information while capturing the evolutionary patterns of time series in a finer granularity. To select more appropriate modeling method based on the characteristics of the dataset, our model unifies the channel-independent and channel-mixing tokenization strategies and uses a series-relation-aware decider to control the strategy choosing process. Extensive experiments on seven real-world datasets show that our model achieves more accurate predictions compared with state-of-the-art methods.
@@ Introduction
Recently, state-space models (SSM) emerge as a promising architecture for sequence modeling(Gu et al. 2021b, a). Among them, a recent study, Mamba(Gu and Dao 2023), has achieved remarkable results in sequence processing tasks such as natural language processing(Lieber et al. 2024), DNA sequence learning, audio waveform processing(Gu and Dao 2023) and computer vision(Zhu et al. 2024). Benefiting from its design of selective scanning, Mamba shows superior performance for long sequence modeling, which makes it potentially suitable for the LTSF task. However, there are limited utilizations of SSM in LTSF currently, which may stem from the inherent challenges in time series analysis tasks as follows:
@@
  • Long-term time series modeling. LTSF deals with larger magnitudes of sequence elements and is often more affected by data non-stationarity, noise and outliers, making it more difficult to capture long-term dependencies. Furthermore, the semantic information density of time series data at time points is lower than other types of sequence data and directly modeling point-wise tokens may introduce redundant noise(Cao et al. 2023). Works such as PatchTST(Nie et al. 2023) and Crossformer(Zhang and Yan 2023) emphasis capturing richer semantic information by dividing time series into patches for tokenization. Compared with point-wise input tokens, patch-wise tokens reduces the number of sequence elements that the model needs to process, which also leads to lower computational complexity. A recent work iTransformer(Liu et al. 2023) uses a simple fully connected layer to map the whole input sequence to hidden states. Although iTransformer gets state-of-the-art (SOTA) performance, it is coarse-grained and is not conducive to capturing fine-grained evolutionary patterns inside the time series. Therefore, we are motivated to model the time series in a patching manner.
@@
  • Channel-independent and channel-mixing strategies. Time series data may have multiple variables which have complex correlations between each other. Due to various internal and external factors, the inter-series dependencies of multivariate time series (MTS) data over a period of time is dynamic. For example, the traffic flow of interconnected roads within a transportation network influence each other in different manners during different periods of time. However, modeling inter-series dependencies using channel-mixing strategies does not always produce SOTA performance. Works such as PatchTST and NHITS(Challu et al. 2023) process the MTS data in a channel-independent way, where each univariate series is input to the model independently to enhance the robustness of the model. A recent study TimeMachine(Ahamed and Cheng 2024) proposes a unified structure to handle channel-mixing and channel-independent situations. However, the boundary for the selection of the two strategies is ambiguous. Directly choosing strategies based on the relationship between the length of the historical observations and the number of variables in the dataset is lack of generality.
@@
To solve the above challenges, we introduce Bi-Mamba4TS, a Bidirectional Mamba for Time Series forecasting. We aim to leverage Mamba’s linear computational complexity and powerful long sequence modeling capabilities. We design a series-relation-aware (SRA) decider based on the Pearson coefficient correlation to automatically choose channel-independent or channel-mixing tokenization strategies for varying datasets. We divide the input sequence into patches and generate patch-wise tokens based on the chosen tokenization strategy. The patch-wise tokens contains richer semantic information and encourage the model to learn the evolutionary patterns of the time series in a finer granularity. Besides, we propose a bidirectional Mamba encoder to model the MTS data more comprehensively. With the above designs, our model can choose the optimal strategy to capture the intra-series dependencies or the inter-series dependencies based on the characteristics of the specific MTS data. The main contributions of this paper are summarized as follows:
@@
  • We propose Bi-Mamba4TS, a long-term multivariate time series forecasting model based on SSM architecture. our model unifies channel-independent and channel-mixing tokenization strategies. Besides, our model divides the time series into patches to capture long-term dependencies in a finer granularity.
  • We design a SRA decider to automatically choose channel-independent or channel-mixing tokenization strategies. The SRA decider provides more objective results based on the Pearson correlation coefficient of different datasets.
@@
  • We conduct extensive experiments on seven widely used real-world datasets. The results show that Bi-Mamba4TS achieves superior performance with varying prediction lengths.
@@
Figure 1: The architecture of Bi-Mamba4TS. The input time series is first divided into patches and then pass through a channel-independent or channel-mixing embedding layer based on the data characteristics of specific datasets. The embeddings are then fed into multiple Bi-Mamba Encoders and get the final output through an MLP projector.
@@
Deep learning models for time series forecasting are initially based on Recurrent Neural Network (RNN) or Convolutional Neural Network (CNN). RNNs process the sequence elements step by step and maintain a hidden state which is updated with each input element. These models are relatively simple and have excellent inference speed. However, the calculated gradient must be passed through all cells one by one, which limits the training speed and leads to forgetting long-term information. CNNs use convolutional kernel to emphasis local information. These models benefit from parallel computing and have faster training speed. However, the convolutional calculating process limits the inference speed and overlook the long-term global information. To solve the issues of these two models, a new model structure is proposed, namely State Space Models (SSM). SSM(Gu et al. 2021b, a) is inspired by the continious system. It is trained in parallel like CNN and inferences fastly like RNN. Some previous works have attempted to use SSM in TSF. SSDNet(Lin et al. 2021) combines the Transformer architecture with SSM to provide probabilistic and interpretable forecasts. SPACETIME(Gu et al. 2021a) proposes a new SSM parameterization based on the companion matrix to enhance the expressivity of the model and introduces a “closed-loop” variation of the companion SSM for long-term time series forecasting.
Recently, a new SSM-based model, Mamba(Gu and Dao 2023), is proposed. It introduces parameterized matrices and a hardware-aware parallel computing algorithm to SSM and achieves superior performance on language modeling, DNA sequence and audio waveform processing tasks. Several Mamba-based derivative models have been developed and used for computer vision tasks(Zhu et al. 2024; Ma et al. 2024) and time series tasks. S/D-Mamba(Wang et al. 2024) explores to use Mamba to capture inter-series dependencies of MTS data. It embeds each univariate time series like iTransformer and feeds the embeddings into Mamba blocks to model the relationships of different time series. However, the tokenization approach may overlook the complex evolutionary patterns inside the time series. TimeMachine(Ahamed and Cheng 2024) proposes a multi-scale quadruple-Mamba architecture to unify the handling of channel-mixing and channel-independence situations. However, the channel-mixing and channel-independent strategies are chosen simply based on the length of historical observations and variable number of different datasets. This approach does not fully consider the characteristics of the MTS data.
@@
The discretized SSM(Gu et al. 2021b) can be trained in parallel in a convolutional operation way and make efficiently inference in a recurrent neural network manner. By introducing HIPPO Matrix(Gu et al. 2020) to the initialization of matrix $\mathbf{A}$, a variant of SSM, namely the structured state space model (S4)(Gu et al. 2021a), makes improvement on the ability to model long-term dependencies.
@@
The overall architecture of Bi-Mamba4TS is shown in Fig. 1. We first calculate the tokenization strategy indicator through the SRA decider. we then divide the input series into patches and generate patch-wise tokens based on the tokenization strategy indicator. The obtained tokens are fed into multiple Bi-Mamba Encoders. We adopt a flatten head and a linear projector to get the final output.
@@
0:  Training set $T=\{t^{1},t^{2},...,t^{M}\}$ of MTS data
@@
4:        $\rho_{i,j}=\rho_{j,i}=\frac{\mathrm{conv}(t^{i},t^{j})}{\sigma_{t^{i}}\cdot\sigma_{t^{j}}}$
@@
Therefore, we design a SRA decider to automatically control the tokenization process of the model. The workflow of the dicision maker is shown in Alg.1. For a specific dataset, we extract the training set data $T$ and calculate the Pearson correlation coefficients of different series $T^{i}$ and $T^{j}$, denoted as $\rho_{i,j}$, where $i$ and $j$ are the indexes of the series ranging from $1$ to $M$ and $i\neq j$. We then use threshold $\lambda$ and 00 to filter out series pairs with positive correlation. Finally, we count the maximum number of relevant series $\rho_{\max}^{\lambda}$ and $\rho_{\max}^{0}$ in the training set and calculate the relation ratio $r=\rho_{\max}^{\lambda}/\rho_{\max}^{0}$. We adopt channel-mixing strategy to generate sequence tokens for datasets with $r\geq 1-\lambda$, otherwise, we adopt channel-independent strategy.
@@
We generalize patch-wise tokens to emphasize capturing local evolutionary patterns of the time series. Specifically, each channel univariate sequence $x^{i}_{1:L}$ is first divided into patches $p^{i}\in\mathbb{R}^{J\times P}$ where $P$ refers to the length of each patch and $J$ is the number of patches. There are two strategies to process the input patches to get patch-wise tokens, namely channel-independent and channel-mixing strategies. For channel-independent strategy, each univariate channel is processed individually and finally concatenated to the embeddings $\mathbb{E}_{ind}\in\mathbb{R}^{M\times J\times D}$, where $D$ is the hidden state dimension. For channel-mixing strategy, we first group patches with the same index of different series and pass each group through the tokenization layer. In this case, the obtained tokens can be represented as $\mathbb{E}_{mix}\in\mathbb{R}^{J\times M\times D}$.
Algorithm 2 Implementation of Bi-Mamba Encoder
0:  Patch-wise token sequence $\mathbb{E}_{x}^{(l)}:(B,S,D)$
0:  Output features for the next encoder layer $\mathbb{E}_{x}^{(l+1)}:(B,S,D)$
@@
3:        $\mathbb{E}_{x,dir}^{(l)}:(B,S,D)\leftarrow\mathrm{Flip}(\mathbb{E}_{x}^{(l)},\mathrm{dim}=1)$
@@
5:        $\mathbb{E}_{x,dir}^{(l)}:(B,S,D)\leftarrow\mathbb{E}_{x}^{(l)}$
@@
7:     $\mathrm{x}:(B,S,E)\leftarrow\mathrm{Linear^{x}}(\mathbb{E}_{x,dir}^{(l)})$
8:     $\mathrm{z}:(B,S,E)\leftarrow\mathrm{Linear^{z}}(\mathbb{E}_{x,dir}^{(l)})$
9:     $\mathrm{x}^{\prime}:(B,S,E)\leftarrow\mathrm{SiLU(Conv1D(x))}$
10:     $\mathbf{A}:(E,N)\leftarrow\mathrm{Parameter}^{\mathbf{A}}$
11:     $\mathbf{B}:(B,S,N)\leftarrow\mathrm{Linear}^{\mathbf{B}}(\mathrm{x}^{\prime})$
12:     $\mathbf{C}:(B,S,N)\leftarrow\mathrm{Linear}^{\mathbf{C}}(\mathrm{x}^{\prime})$
13:     $\Delta:(B,S,E)\leftarrow\log(1+\exp(\mathrm{Linear}^{\Delta}(\mathrm{x}^{\prime}))+\mathrm{Parameter}^{\Delta})$
14:     $\overline{\mathbf{A}},\overline{\mathbf{B}}:(B,S,E,N)\leftarrow\mathrm{discretize(\Delta,\mathbf{A},\mathbf{B})}$
15:     $\mathrm{y}:(B,S,E)\leftarrow\mathrm{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(\mathrm{x}^{\prime})$
16:     $\mathrm{y}^{\prime}:(B,S,E)\leftarrow\mathrm{y}\otimes\mathrm{SiLU(z)}$
17:     $\tilde{\mathbb{E}}_{y,dir}^{(l)}:(B,S,D)\leftarrow\mathrm{Linear^{y}}(\mathrm{y}^{\prime})$
18:     $\tilde{\mathbb{E}}_{y,dir}^{(l)}:(B,S,D)\leftarrow\mathrm{Add\&Norm}(\tilde{\mathbb{E}}_{y,dir}^{(l)},\mathbb{E}_{x,dir}^{(l)})$
19:     $\mathbb{E}_{y,dir}^{(l)}:(B,S,D)\leftarrow\mathrm{FeedForward}(\tilde{\mathbb{E}}_{y,dir}^{(l)})$
20:     $\mathbb{E}_{y,dir}^{(l)}:(B,S,D)\leftarrow\mathrm{Add\&Norm}(\mathbb{E}_{y,dir}^{(l)},\tilde{\mathbb{E}}_{y,dir}^{(l)})$
21:  end for
22:  return $\mathbb{E}_{x}^{(l+1)}:(B,S,D)\leftarrow\mathbb{E}_{y,forward}^{(l)}+\mathbb{E}_{y,backward}^{(l)}$
Bidirectional Mamba Encoder
A original Mamba block is designed to process 1-D sequence on one direction. Considering the diverse representations of features within individual series in different directions, as well as the complexity of inter-series dependencies, we design a bidirectional Mamba structure to comprehensively capture the information of time series.
A Bi-Mamba encoder takes $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{B\times S\times D}$ as input, where $l$ refers to the encoder layer, $B$ and $S$ corresponds to $M$ or $J$ depending on the tokenization strategy. Specifically, if $ts=1$, $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{J\times M\times D}$ and $\mathbb{E}_{x}^{(0)}=\mathbb{E}_{mix}$, otherwise, $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{M\times J\times D}$ and $\mathbb{E}_{x}^{(0)}=\mathbb{E}_{ind}$. There are two Mamba blocks in one Bi-Mamba encoder to model the input sequence from the forward and backward directions respectively, as shown in Fig LABEL:fig:mamba_block(a). We denote the input of these two directions as $\mathbb{E}_{x,dir}^{(l)}$ where $dir\in\{forward,backward\}$. A Mamba block first passes the input $\mathbb{E}_{x,dir}^{(l)}$ through two linear layers to expand the hidden dimension to the size of $E$, as shown in Fig. LABEL:fig:mamba_block(b). The result of one of the branches is passed through a 1-D convolutional layer, a SiLU activation function and a standard SSM block to get the candidate output $\tilde{\mathbb{E}}_{y,dir}^{(l)}$. The result of another branch is passed through a SiLU activation function to serve as a gate. It is multiplied with $\tilde{\mathbb{E}}_{y,dir}^{(l)}$ to get the final output $\mathbb{E}_{y,dir}^{(l)}$. We get $\mathbb{E}_{x}^{(l+1)}=\sum_{dir}^{\{forward,backward\}}\mathcal{F}(\mathbb{E}_{y,dir}^{(l)},\mathbb{E}_{x,dir}^{(l)})$ as the input of the next Bi-Mamba encoder layer. The function $\mathcal{F}$ here represents the combination of residual connection and feed forward neural network layers like those in a Transformer encoder. $\mathbb{E}_{y,forward}^{(l)}$ and $\mathbb{E}_{y,backward}^{(l)}$ are the outputs of the two Mamba blocks with two modeling directions respectively. The detailed implementation is shown in Alg.2.
@@
We choose 7 real-world datasets to evaluate our proposed model: Weather, Traffic, Electricity and 4 ETT datasets(ETTh1, ETTh2, ETTm1, ETTm2). These datasets are widely used, covering multiple fields including weather, transportation and energy management. The statistics of the datasets are shown in Table 1.
Table 1: Statistics of datasets.
@@
Table 2: The information of experimental platform.
@@
GPU model NVIDIA Tesla V100 32GB
@@
Table 3: Experimental results of long-term multivariate time series forecasting task on 7 real-world datasets. The best results are in bold and the second best results are underlined.
Models Bi-Mamba4TS D-Mamba iTransformer PatchTST Crossformer Autoformer DLinear TimesNet CrossGNN WITRAN

### v3

# Bi-Mamba+: Bidirectional Mamba for Time Series Forecasting

@@ Abstract
Long-term time series forecasting (LTSF) provides longer insights into future trends and patterns. Over the past few years, deep learning models especially Transformers have achieved advanced performance in LTSF tasks. However, LTSF faces inherent challenges such as long-term dependencies capturing and sparse semantic characteristics. Recently, a new state space model (SSM) named Mamba is proposed. With the selective capability on input data and the hardware-aware parallel computing algorithm, Mamba has shown great potential in balancing predicting performance and computational efficiency compared to Transformers. To enhance Mamba’s ability to preserve historical information in a longer range, we design a novel Mamba+ block by adding a forget gate inside Mamba to selectively combine the new features with the historical features in a complementary manner. Furthermore, we apply Mamba+ both forward and backward and propose Bi-Mamba+, aiming to promote the model’s ability to capture interactions among time series elements. Additionally, multivariate time series data in different scenarios may exhibit varying emphasis on intra- or inter-series dependencies. Therefore, we propose a series-relation-aware decider that controls the utilization of channel-independent or channel-mixing tokenization strategy for specific datasets. Extensive experiments on 8 real-world datasets show that our model achieves better predictions compared with state-of-the-art methods. Our code is available at https://github.com/Leopold2333/Bi-Mamba+.
@@ Introduction
Recently, state-space models (SSM) emerge as a promising architecture for sequence modeling(Gu et al. 2021a, b). Among them, a recent study, Mamba(Gu and Dao 2023), has achieved remarkable results in sequence processing tasks such as natural language processing(Lieber et al. 2024), DNA sequence learning, audio waveform processing(Gu and Dao 2023) and computer vision(Zhu et al. 2024). Benefiting from its design of selective scanning, Mamba shows superior performance for long sequence modeling, which makes it potentially suitable for the LTSF task. However, there are limited utilizations of SSM in LTSF currently, which may stem from the inherent challenges in time series analysis tasks as follows:
@@
  • Long-term time series modeling. LTSF deals with larger magnitudes of sequence elements and is often more affected by data non-stationarity, noise and outliers, making it more difficult to capture long-term dependencies(Liu et al. 2022). Furthermore, the semantic information density of time series data at time points is lower than other types of sequence data. Directly modeling point-wise tokens may introduce redundant noise(Cao et al. 2023). Works such as PatchTST(Nie et al. 2023) and Crossformer(Zhang and Yan 2023) emphasis capturing richer semantic information by dividing time series into patches for tokenization. Compared with point-wise input tokens, patch-wise tokens reduces the number of sequence elements and leads to lower computational complexity. A recent work iTransformer(Liu et al. 2023) uses a simple fully connected layer to map the whole sequence to hidden states. Although iTransformer gets state-of-the-art (SOTA) performance, it is coarse-grained and is not conducive to capturing fine-grained evolutionary patterns inside the time series. Therefore, we are motivated to model the time series in a patching manner.
@@
  • Emphasis on intra- or inter-series dependencies. Time series data may have multiple variables with complex correlations between each other. For example, the traffic flow of interconnected roads within a transportation network influence each other in different manners during different periods of time. However, capturing inter-series dependencies using channel-mixing strategies does not always produce SOTA performance. Works such as PatchTST and NHITS(Challu et al. 2023) get superior performance by capturing intra-series dependencies only. These models process the multivariate time series (MTS) data in a channel-independent way, where each univariate series is input to the model independently to enhance the robustness of the model. It can be inferred that different datasets shows different emphasis on intra- or inter-sequence dependencies. A recent study TimeMachine(Ahamed and Cheng 2024) proposes a unified structure for channel-independent and channel-mixing tokenization strategies. It aims to handle both intra-series-emphasis and inter-series-emphasis situations. However, the boundary for the selection of tokenization strategies is ambiguous and the statistical characteristics of datasets are overlooked.
@@
To solve the above challenges, we design an improved Mamba block named Mamba+ through adding a forget gate in Mamba to selectively combine the new features with the historical features in a complementary manner, therefore preserving historical information in a longer range. Based on the Mamba+, we propose Bidirectional Mamba+ (Bi-Mamba+) to model the MTS data from both forward and backward, enhancing the model’s robustness and ability to capture interactions between time series elements. In addition, to address the varying emphasis on intra-series evolutionary patterns and inter-series interactions, we design a Series-Relation-Aware (SRA) decider inspired from Spearman coefficient correlation. The SRA decider measures the proportion of highly correlated series pairs in the MTS data to automatically choose channel-independent or channel-mixing tokenization strategies. Furthermore, we divide the input sequence into patches and generate patch-wise tokens based on the chosen tokenization strategy. The patch-wise tokens contain richer semantic information and encourage the model to learn the long-term dependencies of the time series in a finer granularity. The main contributions of this paper are summarized as follows:
  • We propose Bi-Mamba+ for LTSF task. We design an improved Mamba+ block and model the MTS data from both forward and backward to provide more accurate predictions.
@@
  • We design an SRA decider based on the Spearman correlation coefficient to automatically choose channel-independent or channel-mixing tokenization strategies. Furthermore, we divides the time series into patches to capture long-term dependencies in a finer granularity.
@@
  • We conduct extensive experiments on 8 real-world datasets that are widely used. The results show that Bi-Mamba+ achieves superior performance with varying prediction lengths.
@@
Figure 1: The architecture of Bi-Mamba+. The input time series is first divided into patches and then pass through a channel-independent or channel-mixing embedding layer based on the data characteristics of specific datasets. The embeddings are then fed into multiple Bi-Mamba Encoders and get the final output through an MLP projector.
@@
Deep learning models for time series forecasting are initially based on Recurrent Neural Network (RNN) or Convolutional Neural Network (CNN). RNNs process the sequence elements step by step and maintain a hidden state which is updated with each input element. These models are relatively simple and have excellent inference speed. However, the calculated gradient must be passed through all cells one by one, which limits the training speed and leads to forgetting long-term information. CNNs use convolutional kernel to emphasis local information. These models benefit from parallel computing and have faster training speed. However, the convolutional calculating process limits the inference speed and overlook the long-term global information. To solve the issues of these two models, a new model structure is proposed, namely State Space Models (SSM). SSM(Gu et al. 2021a, b) is inspired by the continious system. It is trained in parallel like CNN and inferences fastly like RNN. Some previous works have attempted to use SSM in TSF. SSDNet(Lin et al. 2021) combines the Transformer architecture with SSM to provide probabilistic and interpretable forecasts. SPACETIME(Gu et al. 2021b) proposes a new SSM parameterization based on the companion matrix to enhance the expressivity of the model and introduces a “closed-loop” variation of the companion SSM for long-term time series forecasting.
Recently, a new SSM-based model, Mamba(Gu and Dao 2023), is proposed. It introduces parameterized matrices and a hardware-aware parallel computing algorithm to SSM and achieves superior performance on language modeling, DNA sequence and audio waveform processing tasks. Several Mamba-based derivative models have been developed and used for computer vision tasks(Zhu et al. 2024; Ma et al. 2024) and time series tasks. S-Mamba(Wang et al. 2024) explores to use Mamba to capture inter-series dependencies of MTS data. It embeds each univariate time series like iTransformer and feeds the embeddings into Mamba blocks to model the relationships of different time series. However, the tokenization approach may overlook the complex evolutionary patterns inside the time series. MambaMixer(Behrouz et al. 2024) adjusts the Mamba block to bidirectional and uses two improved blocks to capture inter- and intra-series dependencies simultaneously. However, the gating branch is used to filter new features of both forward and backward directions, which may cause challenges for extracting new features. TimeMachine(Ahamed and Cheng 2024) proposes a multi-scale quadruple-Mamba architecture to unify the handling of channel-mixing and channel-independence situations. However, the channel-mixing and channel-independent strategies are chosen simply based on the length of historical observations and variable number of different datasets. Therefore, the characteristics of the MTS data are not fully considered.
@@
The discretized SSM(Gu et al. 2021a) can be trained in parallel in a convolutional operation way and make efficiently inference in a recurrent neural network manner. By introducing HIPPO Matrix(Gu et al. 2020) to the initialization of matrix $\mathbf{A}$, a variant of SSM, namely the structured state space model (S4)(Gu et al. 2021b), makes improvement on the ability to model long-term dependencies.
@@
The overall architecture of Bi-Mamba+ is shown in Fig. 1. We first calculate the tokenization strategy indicator through the SRA decider. we then divide the input series into patches and generate patch-wise tokens based on the tokenization strategy indicator. The obtained tokens are fed into multiple Bi-Mamba+ encoders. We adopt a flatten head and a linear projector to get the final output.
@@
0:  Training set $T=\{t^{1},t^{2},...,t^{M}\}$ of MTS data, each series has $n$ observations
@@
4:        $\rho_{i,j}=\rho_{j,i}=1-\frac{6\sum_{k=0}^{n}\left(\mathrm{Rank}(t_{k}^{i})-\mathrm{Rank}(t_{k}^{j})\right)^{2}}{n(n^{2}-1)}$
@@
Therefore, we design a SRA decider to automatically control the tokenization process of the model. The workflow of the dicision maker is shown in Alg.1. For a specific dataset, we extract the training set data $T=\{t^{1},t^{2},...,t^{M}\}$ and calculate the Spearman correlation coefficients of different series $t^{i}$ and $t^{j}$, denoted as $\rho_{i,j}$, where $i$ and $j$ are the indexes of the series ranging from $1$ to $M$ and $i\neq j$. We then use threshold $\lambda$ and 00 to filter out series pairs with positive correlation. Finally, we count the maximum number of relevant series $\rho_{\max}^{\lambda}$ and $\rho_{\max}^{0}$ in the training set and calculate the relation ratio $r=\rho_{\max}^{\lambda}/\rho_{\max}^{0}$. We adopt channel-mixing strategy to generate sequence tokens for datasets with $r\geq 1-\lambda$, otherwise, we adopt channel-independent strategy. Spearman coefficient is a nonparametric statistical indicator for evaluating the monotonic relationship between two sequences. It calculates the correlation based on the rank of the sequence elements, making it more suitable for non-linear relationships and more robust to outliers. Given two sequences $t^{i}$ and $t^{j}$, it first sorts the sequence and then define the Spearman coefficient as $\rho_{i,j}=1-\frac{6\sum_{k=0}^{n}\left(\mathrm{Rank}(t_{k}^{i})-\mathrm{Rank}(t_{k}^{j})\right)^{2}}{n(n^{2}-1)}$, where $n$ is the number of observation samples of each sequence and $\mathrm{Rank}(t_{k}^{i})$ represents the rank level of the $k$-th element in the specific time series $t^{i}$.
@@
We generalize patch-wise tokens to emphasize capturing local evolutionary patterns of the time series. Specifically, each channel univariate sequence $x^{i}_{1:L}$ is first divided into patches $p^{i}\in\mathbb{R}^{J\times P}$ where $J$ is the total number of patches and $P$ is the length of each patch. We denote $S$ as the stride among patches and $J=\lceil\frac{L-P}{S}+1\rceil$. For channel-independent strategy, each univariate channel is concatenated to the tokens $\mathbb{E}_{ind}\in\mathbb{R}^{M\times J\times D}$, where $D$ is the hidden state dimension. For channel-mixing strategy, we first group patches with the same index of different series and pass each group through the tokenization layer. The obtained tokens can be represented as $\mathbb{E}_{mix}\in\mathbb{R}^{J\times M\times D}$.
Algorithm 2 Implementation of Mamba+ block
0:  Patch-wise token sequence $\mathbb{E}_{x}:(B,W,D)$
0:  Output features $\mathbb{E}_{y}:(B,W,E)$
1:  $\mathrm{x}:(B,W,E)\leftarrow\mathrm{Linear^{x}}(\mathbb{E}_{x})$
2:  $\mathrm{z}:(B,W,E)\leftarrow\mathrm{Linear^{z}}(\mathbb{E}_{x})$
3:  $\mathrm{x}^{\prime}:(B,W,E)\leftarrow\mathrm{SiLU(Conv1D(x))}$
4:  $\mathbf{A}:(E,N)\leftarrow\mathrm{Parameter}^{\mathbf{A}}$
5:  $\mathbf{B}:(B,W,N)\leftarrow\mathrm{Linear}^{\mathbf{B}}(\mathrm{x}^{\prime})$
6:  $\mathbf{C}:(B,W,N)\leftarrow\mathrm{Linear}^{\mathbf{C}}(\mathrm{x}^{\prime})$
7:  $\Delta:(B,W,E)\leftarrow\log(1+\exp(\mathrm{Linear}^{\Delta}(\mathrm{x}^{\prime}))+\mathrm{Parameter}^{\Delta})$
8:  $\overline{\mathbf{A}},\overline{\mathbf{B}}:(B,W,E,N)\leftarrow\mathrm{discretize(\Delta,\mathbf{A},\mathbf{B})}$
9:  $\mathrm{y}:(B,W,E)\leftarrow\mathrm{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(\mathrm{x}^{\prime})$
10:  $\mathrm{y}^{\prime}:(B,W,E)\leftarrow\mathrm{y}\otimes\mathrm{SiLU(z)}+\mathrm{x}^{\prime}\otimes(1-\sigma(\mathrm{z}))$
11:  $\mathbb{E}_{y}:(B,W,D)\leftarrow\mathrm{Linear^{y^{\prime}}(y^{\prime})}$
12:  return $\mathbb{E}_{y}$
Mamba+ Block
An original Mamba block uses two branches to process the input features. One of the branches, denoted as $b_{1}$, passes the input features through a 1-D convolutional module and a SSM block. The other branch, denoted as $b_{2}$, feeds the input features into a SiLU activation function to serve as a gate. Although the HIPPO matrix(Gu et al. 2021b) embedded in the SSM block can retain a fairly long-term historical information, the obtained result is filtered directly through the gate of another branch, resulting in the tendency to prioritize proximal information(Wang et al. 2024). Therefore, we propose an improved Mamba+ block specifically designed for LTSF. Particularly, we add a forget gate calculated by $gate_{f}=1-gate_{b_{2}}$, where $gate_{f}$ represents the forget gate and $gate_{b_{2}}$ is the result of sigmoid function in $b_{2}$. The output of the 1-D convolutional module $x^{\prime}$ is multiplied with $gate_{f}$ and then added to the filtered result of SSM. $gate_{f}$ and $gate_{b_{2}}$ selectively combine the added new features with the forgotten historical features in a complementary manner. Compared to directly filtering the obtained new features, Mamba+ emphasis preserving historical information in a longer range. The detailed flow of Mamba+ is shown in Alg.2. The reprogrammed forward and backward propagation process of the hardware-aware parallel computing algorithm is shown in Appendix.
Algorithm 3 Implementation of Bi-Mamba+ Encoder
0:  Patch-wise token sequence $\mathbb{E}_{x}^{(l)}:(B,W,D)$
0:  Output features for the next encoder layer $\mathbb{E}_{x}^{(l+1)}:(B,W,D)$
@@
3:        $\mathbb{E}_{x,dir}^{(l)}:(B,W,D)\leftarrow\mathrm{Flip}(\mathbb{E}_{x}^{(l)},\mathrm{dim}=1)$
@@
5:        $\mathbb{E}_{x,dir}^{(l)}:(B,W,D)\leftarrow\mathbb{E}_{x}^{(l)}$
@@
7:     $\tilde{\mathbb{E}}_{y,dir}^{(l)}:(B,W,D)\leftarrow\mathrm{TSMamba}(\mathbb{E}_{x,dir}^{(l)})$
8:     $\tilde{\mathbb{E}}_{y,dir}^{(l)}:(B,W,D)\leftarrow\mathrm{Add\&Norm}(\tilde{\mathbb{E}}_{y,dir}^{(l)},\mathbb{E}_{x,dir}^{(l)})$
9:  end for
10:  $\tilde{\mathbb{E}}_{y}^{(l+1)}:(B,W,D)\leftarrow\mathbb{E}_{y,forward}^{(l)}+\mathbb{E}_{y,backward}^{(l)}$
11:  $\tilde{\mathbb{E}}_{x}^{(l+1)}:(B,W,D)\leftarrow\mathrm{FeedForward}(\tilde{\mathbb{E}}_{y}^{(l+1)})$
12:  $\mathbb{E}_{x}^{(l+1)}:(B,W,D)\leftarrow\mathrm{Add\&Norm}(\tilde{\mathbb{E}}_{x}^{(l+1)},\tilde{\mathbb{E}}_{y}^{(l+1)})$
13:  return $\mathbb{E}_{x}^{(l+1)}$
Bidirectional Mamba+ Encoder
An original Mamba block is designed to process 1-D sequence on one direction. Considering the rich evolutionary patterns of time series in different directions as well as the complexity of inter-series dependencies, we design a bidirectional Mamba+ structure to comprehensively model the MTS data.
A Bi-Mamba+ encoder takes $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{B\times W\times D}$ as input, where $l$ refers to the encoder layer, $B$ and $W$ corresponds to $M$ or $J$ depending on the tokenization strategy. Specifically, if $ts=1$, $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{J\times M\times D}$ and $\mathbb{E}_{x}^{(0)}=\mathbb{E}_{mix}$, otherwise, $\mathbb{E}_{x}^{(l)}\in\mathbb{R}^{M\times J\times D}$ and $\mathbb{E}_{x}^{(0)}=\mathbb{E}_{ind}$. There are two Mamba+ blocks in one Bi-Mamba+ encoder to model the input sequence from the forward and backward directions respectively, as shown in Fig LABEL:fig:mamba_block(a). We denote the input of these two directions as $\mathbb{E}_{x,dir}^{(l)}$ where $dir\in\{forward,backward\}$. We get $\mathbb{E}_{x}^{(l+1)}=\sum_{dir}^{\{forward,backward\}}\mathcal{F}(\mathbb{E}_{y,dir}^{(l)},\mathbb{E}_{x,dir}^{(l)})$ as the input of the next Bi-Mamba+ encoder layer. The function $\mathcal{F}$ here represents the combination of residual connection and feed forward neural network layers like those in a Transformer encoder. $\mathbb{E}_{y,forward}^{(l)}$ and $\mathbb{E}_{y,backward}^{(l)}$ are the outputs of the two Mamba+ blocks with two modeling directions respectively. The detailed implementation is shown in Alg.3.
@@
We choose 8 real-world datasets to evaluate our proposed model: Weather, Traffic, Electricity, Solar and 4 ETT datasets(ETTh1, ETTh2, ETTm1, ETTm2). These datasets are widely used, covering multiple fields including weather, transportation and energy management. The statistics of the datasets are shown in Table 1.
Table 1: Statistics of all datasets.
@@
Solar 137 10min 52179
Table 2: The information of the experimental platform.
@@
GPU model NVIDIA Tesla V100 32GB (×2)($\times$ 2)( × 2 )
@@
Table 3: Experimental results of long-term multivariate time series forecasting task on 8 real-world datasets. The best results are in bold and the second best results are underlined.
Models Bi-Mamba+ S-Mamba iTransformer PatchTST Crossformer Autoformer DLinear TimesNet CrossGNN WITRAN